### v4

@@ 3   DID-IV in multiple time periods
We now extend the $2\times 2$ DID-IV design to multiple period settings with the staggered adoption of the instrument across units. In reality, many studies exploit variation in the timing of policy adoption across units as an instrument for treatment in more than two periods (e.g. Black et al. (2005), Bhuller et al. (2013), Lundborg et al. (2014), and Meghir et al. (2018)). We refer to the underlying identification strategy as a staggered DID-IV design, and establish the target parameter and identifying assumptions.
@@ 3.1   Set up
We introduce the notation we use throughout Section 3 to Section 5. We consider a panel data setting with $T$ periods and $N$ units. For each $i\in\{1,\dots N\}$ and $t\in\{1,\dots,T\}$, let $Y_{i,t}$ denote the outcome, $D_{i,t}\in\{0,1\}$ denote the treatment status, and $Z_{i,t}\in\{0,1\}$ denote the instrument status. Let $D_{i}=(D_{i,1},\dots,D_{i,T})$ and $Z_{i}=(Z_{i,1},\dots,Z_{i,T})$ denote the path of the treatment and the instrument for unit $i$, respectively. Throughout Section 3 to Section 5, we assume that $\{Y_{i,t},D_{i,t},Z_{i,t}\}_{t=1}^{T}$ are independent and identically distributed (i.i.d).
@@
Assumption 7 requires that once units start getting exposed to the instrument, units remain exposed to that instrument. In the DID literature, several studies make a similar assumption for the treatment and call it the ”staggered treatment adoption” (e.g. Athey and Imbens (2022); Callaway and Sant’Anna (2021); and Sun and Abraham (2021)).
Given Assumption 7, one can uniquely characterize one’s instrument path by the initial exposure date of the instrument, denoted as $E_{i}=\min\{t:Z_{i,t}=1\}$. If unit $i$ is not exposed to the instrument for all time periods, we define $E_{i}=\infty$. Based on the initial exposure period $E_{i}$, one can uniquely partition units into mutually exclusive and exhaustive cohorts $e$ for $e\in\{1,2,\dots,T,\infty\}$: all the units in cohort $e$ start being exposed to the instrument in period $E_{i}=e$.
@@
Assumption 7 allows us to rewrite $D_{i,t}(z)$ by the initial exposure date $E_{i}=e$. Let $D_{i,t}^{e}$ denote the potential treatment status in period $t$ if unit $i$ is first exposed to the instrument in period $e$. Let $D_{i,t}^{\infty}$ denote the potential treatment status in period $t$ if unit $i$ is never exposed to the instrument. We call $D_{i,t}^{\infty}$ the ”never exposed treatment”. Since the adoption date of the instrument uniquely pins down one’s instrument path, we can write the observed treatment status $D_{i,t}$ for unit $i$ in period $t$ as
@@
We define $D_{i,t}-D_{i,t}^{\infty}$ to be the effect of an instrument on treatment for unit $i$ in period $t$, which is the difference between the observed treatment status $D_{i,t}$ to the never exposed treatment status $D_{i,t}^{\infty}$. We refer to $D_{i,t}-D_{i,t}^{\infty}$ as the individual exposed effect in the first stage. In the DID literature, Callaway and Sant’Anna (2021) and Sun and Abraham (2021) define the effect of a treatment on an outcome in the same fashion.
Next, we introduce the group variable that describes the type of unit $i$ in period $t$, based on the reaction of potential treatment choices in period $t$ to the instrument path $z$. Let $G_{i,e,t}\equiv(D_{i,t}^{\infty},D_{i,t}^{e})$ ($t\geq e$) be the group variable in period $t$ for unit $i$ and the initial exposure date $e$. The first element $D_{i,t}^{\infty}$ represents the treatment status in period $t$ if unit $i$ is never exposed to the instrument $E_{i}=\infty$ and the second element $D_{i,t}^{e}$ represents the treatment status in period $t$ if unit $i$ is first exposed to the instrument at $E_{i}=e$. Following the terminology in section 2.1, we define $G_{i,e,t}=(0,0)\equiv NT_{e,t}$ as the never-takers, $G_{i,e,t}=(1,1)\equiv AT_{e,t}$ as the always-takers, $G_{i,e,t}=(0,1)\equiv CM_{e,t}$ as the compliers and $G_{i,e,t}=(1,0)\equiv DF_{e,t}$ as the defiers in period $t$ and the initial exposure date $e$.
@@
Assumption 8 (No carryover assumption).
@@
Henceforth, we keep Assumption 7 and 8. In the next section, we define the target parameter in staggered DID-IV designs.
@@ 3.2   Target parameter in staggered DID-IV designs
Each CLATT is a natural generalization of the LATET in Subsection 2.2 and suitable for the setting of the staggered instrument adoption. This parameter measures the treatment effects at a given relative period $l$ from the initial exposure date $E_{i}=e$, for those who belong to cohort $e$, and are the compliers $CM_{e,e+l}$ who are induced to treatment by instrument in period $e+l$. Each CLATT can potentially vary across cohorts and over time because it depends on cohort $e$, relative period $l$, and the compliers $CM_{e,e+l}$. Note that the composition of the compliers can overlap over the post-exposure periods in cohort $e$. For instance, some units in cohort $e$ may belong to $CM_{e,e+l}$ for all the relative periods ($l\geq e$): they are affected by the instrument for all post-exposure periods.
This parameter would be policy-relevant if the instrument is based on the policy change of interest to the researcher. For example, suppose that the treatment represents participation in a job training program, the outcome represents real earnings, and the instrument represents a policy shock that encourages people to receive the treatment. In this setting, if the instrument is an important policy lever for the researchers (e.g., nudge), each CLATT is of particular interest; the share of compliers in each relative period after the policy shock represents the dynamic effect of that policy change on the take-up rate of the program, and the corresponding CLATT can be interpreted as estimating the effect of participation in job training on real earnings among those affected.
@@ 3.3   Identification assumptions in staggered DID-IV designs
Since the initial exposure date $E_{i}$ completely characterizes the instrument path, we can write the potential outcomes for cohort $e$ and cohort $\infty$ as $Y_{i,t}(D_{i,t}^{e})$ and $Y_{i,t}(D_{i,t}^{\infty})$, respectively. The potential outcome $Y_{i,t}(D_{i,t}^{e})$ represents the outcome status in period $t$ if unit $i$ is first exposed to the instrument in period $e$ and the potential outcome $Y_{i,t}(D_{i,t}^{\infty})$ represents the outcome status in period $t$ if unit $i$ is never exposed. We refer to $Y_{i,t}(D_{i,t}^{\infty})$ as the ”never exposed outcome”.
@@
Assumption 12 is a parallel trends assumption in the treatment in multiple periods and cohorts. It requires that the trends of the treatment across cohorts would have followed the same path, on average, if there was no exposure to the instrument. Assumption 12 is analogous to that of Callaway and Sant’Anna (2021) and Sun and Abraham (2021) in DID designs; these studies impose the same type of parallel trends assumption on untreated outcomes in settings of multiple periods and cohorts.
@@
Assumption 13 is a parallel trends assumption in the outcome with settings of multiple periods and cohorts. It requires that the expectation of the never exposed outcome across cohorts would have followed the same evolution if the instrument had not been assigned. From the discussion in Subsection 2.4, this assumption requires the same expected time gain across cohorts and over time: the effects of time on the outcome through treatment are the same on average across cohorts and over time.
@@ 4   Extensions
This section contains extensions to non-binary, ordered treatments, triple DID-IV designs, and repeated cross sections. It includes the case in which the treatment path was introduced in potential outcomes in $2\times 2$ DID-IV designs. For more details and proofs, see Appendix B.
@@
Up to now, we have considered only the case of a binary treatment. However the same idea can be applied when treatment takes a finite number of ordered values; that is, we have $D_{i,t}\in\{0,1,\dots,J\}$. When we have only two periods and treatment is non-binary, our target parameter is the average causal response on the treated (ACRT) defined below.
@@
$ACRT\equiv\sum_{j=1}^{J}w_{j}\cdot E[Y_{1}(j)-Y_{1}(j-1)|D_{i,1}((0,1))\geq j>D_{i,1}((0,0)),E_{i}=1].$
@@
$w_{j}=\frac{Pr(D_{1}((0,1))\geq j>D_{1}((0,0))|E_{i}=1)}{\sum_{j=1}^{J}Pr(D_{1}((0,1))\geq j>D_{1}((0,0))|E_{i}=1)}.$
The ACRT is a weighted average of the effects of an increase of one unit in the treatment on the outcome, for those who belong to an exposed group and are induced to increase the treatment in period $t=1$ by instrument. The ACRT is similar to the average causal response (ACR) considered in Angrist and Imbens (1995), with the difference that each weight $w_{j}$ and the associated causal parameter in this parameter are conditional on $E_{i}=1$.
Theorem 3 below shows that if we have a non-binary, ordered treatment, the Wald-DID estimand is equal to the ACRT under the same assumptions in Theorem 1.
Theorem 3.
If Assumptions 1-6 hold, the Wald-DID estimand $W_{DID}$ is equal to the ACRT.
$W_{DID}$ $=\sum_{j=1}^{J}w_{j}\cdot E[Y_{1}(j)-Y_{1}(j-1)|D_{i,1}((0,1))\geq j>D_{i,1}((0,0)),E_{i}=1]$
$=ACRT.$
Proof.
See Appendix B.1. ∎
@@
In some applications, researchers have no access to panel data and use repeated cross section data [^6] It also includes the case that researchers use the cross section data, and exploit a policy shock across cohorts as an instrument for the treatment as in Duflo (2001).. In Appendix B.2, we consider repeated cross section data settings and present the identification assumptions in DID-IV designs.
@@
In Subsection 2.1, we impose Assumption 1 (no carryover assumption) on potential outcomes $Y_{i,t}(d,z)$ as we consider the non-staggered adoption of the treatment. We can extend our results to the case where we introduce the treatment path in potential outcomes with binary treatment. In this case, our target parameter changes slightly with some complex notation.
Definition.
$LATET^{\prime}\equiv w_{0}\cdot\Delta_{0}+w_{1}\cdot\Delta_{1}$
where $\Delta_{0}$ and $\Delta_{1}$ are:
The weights $w_{0}$ and $w_{1}$ are:
Appendix B.3 shows that the similar assumptions to Assumptions 1-6 in Subsection 2.3 are sufficient for the Wald-DID to capture the LATET’.
@@ 5   Estimation and inference
When researchers leverage variation in policy adoption timing across units as an instrument for treatment and implicitly rely on a staggered DID-IV design, they usually implement this design via TWFEIV regressions (e.g., Johnson and Jackson (2019), Lundborg et al. (2014), Black et al. (2005), Akerman et al. (2015), and Bhuller et al. (2013)).
The companion paper (Miyaji (2024)), however, shows that in more than two periods, the TWFEIV estimand potentially fails to summarize the treatment effects under staggered DID-IV designs in the presence of heterogeneous treatment effects. Specifically, it shows that under staggered DID-IV designs, the TWFEIV estimand is a weighted average of all possible CLATTs, but some weights can be negative if the effect of the instrument on the treatment or the outcome is not stable over time. [^7] See also the discussions in de Chaisemartin and D’Haultfœuille (2020), who decompose the numerator and denominator in the TWFEIV estimand separately, and point out the issue of interpreting this estimand causally.
In this section, we propose an alternative estimation method in staggered DID-IV designs that is robust to treatment effect heterogeneity. First, we propose a simple regression-based method for estimating each CLATT. Following Callaway and Sant’Anna (2021), we then propose a weighting scheme to construct the summary causal parameters from each CLATT. At the end of this section, we also discuss the pre-trend tests for checking the plausibility of parallel trends assumption in the treatment and the outcome in DID-IV designs.
@@ 5.1   Stacked two stage least squares regression
We propose a regression-based method to consistently estimate our target parameter in staggered DID-IV designs. If we have a binary treatment, the target parameter is the $CLATT_{e,e+l}$ for each cohort $e$ and a relative period $l\geq 0$:
$CLATT_{e,e+l}=E[Y_{i,e+l}(1)-Y_{i,e+l}(0)|E_{i}=e,CM_{e,e+l}].$
Our estimation method consists of two steps.
@@
We create the data sets for each $CLATT_{e,e+l}$. Each data set includes the units of time $t=e-1$ and $t=e+l$, who are either in cohort $e\in\{2,\dots,T\}$ or in the set of some unexposed cohorts, $U$ ($e\notin U$).
Researchers should carefully specify the choice set $U$ in practice. If a never exposed cohort $E_{i}=\infty$ exists, we can set $U=\{\infty\}$ and estimate $CLATT_{e,e+l}$ for all cohorts ($e\in\{2,\dots,T\}$) and all relative time periods ($l\geq 0$). If no never exposed cohorts exist, we can set $U=\max\{E_{i}\}$, the last exposed cohort, and can estimate $CLATT_{e,e+l}$ for cohort $e\in\{2,\dots,\max\{E_{i}\}-1\}$ and relative time period $l\in\{0,\dots,\max\{E_{i}\}-1-e\}$ because every unit will be exposed to the instrument after time $t=\max\{E_{i}\}$. Note that the units in an already exposed cohort $e=1$ are not included in any data sets because the pre-exposed period $e-1$ should be included for estimating each $CLATT_{e,e+l}$.
@@
$Y_{i,t}=\beta^{e,l}_{0}+\beta^{e,l}_{i,.}\mathbf{1}\{E_{i}=e\}+\beta^{e,l}_{,.t}\mathbf{1}\{T_{i}=e+l\}+\beta^{e,l}_{IV}D_{i,t}+\epsilon^{e,l}_{i,t}.$
The first stage regression is:
$D_{i,t}=\pi^{e,l}_{0}+\pi^{e,l}_{i,.}\mathbf{1}\{E_{i}=e\}+\pi^{e,l}_{,.t}\mathbf{1}\{T_{i}=e+l\}+\pi^{e,l}(\mathbf{1}\{E_{i}=e\}\cdot\mathbf{1}\{T_{i}=e+l\})+\eta^{e,l}_{i,t},$
@@
Here, $E_{N}[\cdot]$ is the sample analog of the conditional expectation. Note that the IV estimator $\widehat{CLATT}_{e,e+l}$ is the Wald-DID estimator for each $CLATT_{e,e+l}$, where the choice set $U$ is an unexposed group and time $t=e-1$ is a pre-exposed period. From the estimation procedure, we call this a stacked two-stage least squares (STS) estimator.
Our STS estimator is related to the DID estimators in staggered DID designs proposed by Callaway and Sant’Anna (2021) and Sun and Abraham (2021). The DID estimator of the treatment and the outcome in our STS estimator corresponds to that of Sun and Abraham (2021), and coincides with that of Callaway and Sant’Anna (2021) for the case when no covariates exist and never treated units are used as a control group. Their proposed methods avoid the issue of TWFE estimators in staggered DID designs, and our STS estimator avoids the issue of TWFEIV estimators in staggered DID-IV designs.
@@
Suppose Assumptions 7-13 hold. Then, the STS estimator $\widehat{CLATT}_{e,e+l}$ is consistent and asymptotically normal.
@@
Our estimation procedure is identical if we have a non-binary, ordered treatment or use repeated cross section data. Appendix C presents the influence function of the SLS estimator in repeated cross section settings.
In triple DID-IV designs, one can replace the IV regression in step $2$ with the following one ($g_{i}=A,B$):
$Y_{i,t}$ $=\beta^{e,l}_{0}+\beta^{e,l}_{1}\mathbf{1}_{E_{i}=e}+\beta^{e,l}_{2}\mathbf{1}_{T_{i}=e+l}+\beta^{e,l}_{3}\mathbf{1}_{g_{i}=A}+\beta^{e,l}_{4}\mathbf{1}_{E_{i}=e,T_{i}=e+l}+\beta^{e,l}_{5}\mathbf{1}_{E_{i}=e,g_{i}=A}+\beta^{e,l}_{6}\mathbf{1}_{T_{i}=e+l,g_{i}=A}$
$+\beta^{e,l}_{IV}D_{i,t}+\epsilon^{e,l}_{i,t}.$
The first stage regression is:
$D_{i,t}$ $=\pi^{e,l}_{0}+\pi^{e,l}_{1}\mathbf{1}_{E_{i}=e}+\pi^{e,l}_{2}\mathbf{1}_{T_{i}=e+l}+\pi^{e,l}_{3}\mathbf{1}_{g_{i}=A}+\pi^{e,l}_{4}\mathbf{1}_{E_{i}=e,T_{i}=e+l}+\pi^{e,l}_{5}\mathbf{1}_{E_{i}=e,g_{i}=A}+\pi^{e,l}_{6}\mathbf{1}_{T_{i}=e+l,g_{i}=A}$
$+\pi^{e,l}_{7}\mathbf{1}_{E_{i}=e,T_{i}=e+l,g_{i}=A}+\eta^{e,l}_{i,t},$
where $\mathbf{1}_{A}$ is the indicator function and takes one if $A$ is true. Appendix C presents the influence function of the triple DID-IV estimator $\beta_{IV}^{e,l}$.

### v5

@@ 3   DID-IV in multiple time periods
We now extend the $2\times 2$ DID-IV design to multiple period settings with the staggered adoption of the instrument across units (Black et al. (2005); Bhuller et al. (2013) Lundborg et al. (2014); and Meghir et al. (2018)). We call it a staggered DID-IV design, and establish the target parameter and identifying assumptions.
@@ 3.1   Set up
We introduce the notation we use throughout Section 3 to Section 5. We consider a panel data setting with $T$ periods and $N$ units. For each $i\in\{1,\dots N\}$ and $t\in\{1,\dots,T\}$, let $Y_{i,t}$ denote the outcome, $D_{i,t}\in\{0,1\}$ denote the treatment status, and $Z_{i,t}\in\{0,1\}$ denote the instrument status. Let $D_{i}=(D_{i,1},\dots,D_{i,T})$ and $Z_{i}=(Z_{i,1},\dots,Z_{i,T})$ denote the path of the treatment and the instrument for unit $i$, respectively. Throughout Section 3 to Section 5, we assume that $\{Y_{i,t},D_{i,t},Z_{i,t}\}_{t=1}^{T}$ are i.i.d.
@@
Assumption 8 requires that once units start getting exposed to the instrument, units remain exposed to that instrument. In the DID literature, several studies make a similar assumption for the treatment and call it the “staggered treatment adoption” (e.g. Athey and Imbens (2022); Callaway and Sant’Anna (2021); and Sun and Abraham (2021)).
Given Assumption 8, one can uniquely characterize one’s instrument path by the initial exposure date of the instrument, denoted as $E_{i}=\min\{t:Z_{i,t}=1\}$. If unit $i$ is not exposed to the instrument for all time periods, we define $E_{i}=\infty$. Based on the initial exposure period $E_{i}$, one can uniquely partition units into mutually exclusive and exhaustive cohorts $e$ for $e\in\{1,2,\dots,T,\infty\}$.
@@
Assumption 8 allows us to rewrite $D_{i,t}(z)$ by the initial exposure date $E_{i}=e$. Let $D_{i,t}^{e}$ denote the potential treatment status in period $t$ if unit $i$ is first exposed to the instrument in period $e$. Let $D_{i,t}^{\infty}$ denote the potential treatment status in period $t$ if unit $i$ is never exposed to the instrument. We call $D_{i,t}^{\infty}$ the “never exposed treatment”. Since the adoption date of the instrument uniquely pins down one’s instrument path, we can write the observed treatment status $D_{i,t}$ for unit $i$ in period $t$ as
@@
We define the effect of an instrument on treatment for unit $i$ in period $t$ as the difference between the observed treatment status to the never exposed treatment status: $D_{i,t}-D_{i,t}^{\infty}$. We refer to $D_{i,t}-D_{i,t}^{\infty}$ as the individual exposed effect in the first stage. [^15] In the DID literature, Callaway and Sant’Anna (2021) and Sun and Abraham (2021) define the effect of a treatment on an outcome in the same fashion.
Next, we introduce the group variable that describes the type of unit $i$ in period $t$, based on the reaction of potential treatment choices in period $t$ to the instrument path $z$. Let $G_{i,e,t}\equiv(D_{i,t}^{\infty},D_{i,t}^{e})$ ($t\geq e$) be the group variable in period $t$ for unit $i$ and the initial exposure date $e$. Following the terminology in section 2.1, we define $G_{i,e,t}=(0,0)\equiv NT_{e,t}$ as the never-takers, $G_{i,e,t}=(1,1)\equiv AT_{e,t}$ as the always-takers, $G_{i,e,t}=(0,1)\equiv CM_{e,t}$ as the compliers and $G_{i,e,t}=(1,0)\equiv DF_{e,t}$ as the defiers in period $t$ and the initial exposure date $e$.
@@
Assumption 9 (No carryover assumption in multiple time periods).
@@
Henceforth, we keep Assumption 8 and Assumption 9. In the next section, we define the target parameter in staggered DID-IV designs.
@@ 3.2   Target parameter in staggered DID-IV designs
Each CLATT is a natural generalization of the LATET in Subsection 2.2 and suitable for the setting of the staggered instrument adoption. This parameter measures the treatment effects at a given relative period $l$ from the initial exposure date $E_{i}=e$, for those who belong to cohort $e$, and are the compliers $CM_{e,e+l}$ who are induced to treatment by instrument in period $e+l$. Each CLATT can potentially vary across cohorts and over time because it depends on cohort $e$, relative period $l$, and the compliers $CM_{e,e+l}$. [^16] Note that the composition of the compliers can overlap over the post-exposure periods in cohort $e$. For instance, some units in cohort $e$ may belong to $CM_{e,e+l}$ for all the relative periods ($l\geq e$): they are affected by the instrument for all post-exposure periods.
This parameter would be policy-relevant if the instrument is based on the policy change of interest to the researcher. For example, in a job training program setting considered in Subsection 2.5, each CLATT would be of particular interest if the instrument $Z_{i,t}$ represents an important policy lever for the researchers (e.g., nudge); the share of compliers in each relative period after the policy shock represents the dynamic effect of that policy change on the take-up rate of the program, and the corresponding CLATT can be interpreted as estimating the effect of participation in job training on real earnings among those affected.
@@ 3.3   Identification assumptions in staggered DID-IV designs
Since the initial exposure date $E_{i}$ completely characterizes the instrument path, we can write the potential outcomes for cohort $e$ and cohort $\infty$ as $Y_{i,t}(D_{i,t}^{e})$ and $Y_{i,t}(D_{i,t}^{\infty})$, respectively. The potential outcome $Y_{i,t}(D_{i,t}^{e})$ represents the outcome status in period $t$ if unit $i$ is first exposed to the instrument in period $e$ and the potential outcome $Y_{i,t}(D_{i,t}^{\infty})$ represents the outcome status in period $t$ if unit $i$ is never exposed. We refer to $Y_{i,t}(D_{i,t}^{\infty})$ as the “never exposed outcome”.
@@
Assumption 13 is a parallel trends assumption in the treatment in multiple periods and cohorts. It requires that the trends of the treatment across cohorts would have followed the same path, on average, if there was no exposure to the instrument. Assumption 13 is analogous to that of Callaway and Sant’Anna (2021) and Sun and Abraham (2021) in DID designs; these studies impose the same type of parallel trends assumption on untreated potential outcomes.
@@
Assumption 14 is a parallel trends assumption in the outcome with settings of multiple periods and cohorts. It requires that the expectation of the never exposed outcome across cohorts would have followed the same evolution if the instrument had not been assigned. From the discussion in Subsection 2.4, this assumption can be interpreted as requiring the same expected time gain across cohorts and over time; that is, the effects of time on the outcome through treatment are the same on average across cohorts and over time.
Assumption 15 (Relevance condition in multiple time periods).
$\forall e\in\mathcal{S}(E_{i}),\hskip 5.69054pt\forall l\geq 0,\hskip 5.69054ptE[D_{i,e+l}-D^{\infty}_{i,e+l}|E_{i}=e]>0.$
Assumption 15 is a relevance condition in multiple period settings. With a binary treatment, Assumption 11 and 15 implies that the fraction of the compliers should be non-negative for all $e$ and $l\geq 0$:
$E[D_{i,e+l}-D^{\infty}_{i,e+l}|E_{i}=e]=Pr(CM_{e,l}|E_{i}=e)>0.$
This condition ensures that each $CLATT_{e,l}$ is well defined.
@@ 4   Extensions
This section contains extensions to non-binary, ordered treatments and repeated cross sections. It also includes the case that the treatment path is introduced in potential outcomes in $2\times 2$ DID-IV designs. For more details and proofs, see online Appendix Section A.
@@
Up to now, we have considered only the case of a binary treatment. However the same idea can be applied when treatment takes a finite number of ordered values: $D_{i,t}\in\{0,1,\dots,J\}$. When only two periods exist and treatment is non-binary, our target parameter is the average causal response on the treated (ACRT) defined below.
@@
$ACRT\equiv\sum_{j=1}^{J}w_{j}\cdot E[Y_{i,1}(j)-Y_{i,1}(j-1)|D_{i,1}((0,1))\geq j>D_{i,1}((0,0)),E_{i}=1].$
@@
$w_{j}=\frac{Pr(D_{i,1}((0,1))\geq j>D_{i,1}((0,0))|E_{i}=1)}{\sum_{j=1}^{J}Pr(D_{i,1}((0,1))\geq j>D_{i,1}((0,0))|E_{i}=1)}.$
The ACRT is a weighted average of the effect of one unit increase in the treatment on the outcome, for those who belong to an exposed group and are induced to increase the treatment in period $t=1$ by instrument. The ACRT is similar to the average causal response (ACR) considered in Angrist and Imbens (1995), with the difference that each weight $w_{j}$ and the associated causal parameter in the ACRT are conditional on $E_{i}=1$.
Online Appendix Section A.1 shows that if we have a non-binary, ordered treatment, the Wald-DID estimand captures the ACRT under the same assumptions in Theorem 1.
@@
In some applications, researchers have only access to repeated cross section data. [^17] It also includes the case that researchers use the cross section data, and exploit a policy shock across cohorts as an instrument for treatment as in Duflo (2001). Online Appendix Section A.2 presents the identification assumptions in DID-IV designs under repeated cross section settings.
@@
In Subsection 2.1, we impose Assumption 1 (no carryover assumption) on potential outcomes $Y_{i,t}(d,z)$, as we consider the non-staggered adoption of the treatment. We can extend our results to the case that we introduce the treatment path in potential outcomes with binary treatment. Online Appendix Section A.3 shows that the similar assumptions to Assumptions 1-6 are sufficient for the Wald-DID to capture the LATET-like parameter.
@@ 5   Estimation and inference
In this section, we propose a credible estimation method in staggered DID-IV designs that is robust to treatment effect heterogeneity. First, we propose a simple regression-based method for estimating each CLATT. Following Callaway and Sant’Anna (2021), we then propose a weighting scheme to construct the summary causal parameters from each CLATT. We also discuss the pre-trends tests for checking the plausibility of parallel trends assumptions in DID-IV designs.
Remark 1.
In practice, when researchers implicitly rely on a staggered DID-IV design, they commonly implement this design via two-way fixed instrumental variable (TWFEIV) regressions (e.g., Johnson and Jackson (2019), Lundborg et al. (2014), Black et al. (2005), Akerman et al. (2015), and Bhuller et al. (2013)):
$Y_{i,t}=\phi_{i.}+\lambda_{t.}+\beta_{IV}D_{i,t}+v_{i,t},$
$D_{i,t}=\gamma_{i.}+\zeta_{t.}+\pi Z_{i,t}+\eta_{i,t}.$
In companion paper (Miyaji (2024)), however, we show that in more than two periods, the TWFEIV estimand potentially fails to summarize the treatment effects in the presence of heterogeneous treatment effects. Specifically, we show that the TWFEIV estimand is a weighted average of all possible CLATTs under staggered DID-IV designs, but some weights can be negative if the effect of the instrument on the treatment or the outcome is not stable over time.
@@ 5.1   Stacked two stage least squares regression
We propose a regression-based method to consistently estimate our target parameter in staggered DID-IV designs. Our estimation method consists of two steps.
@@
We create the data sets for each $CLATT_{e,e+l}$. Each data set includes the units of time $t=e-1$ and $t=e+l$, who are either in cohort $e\in\{2,\dots,T\}$ or in the set of some unexposed cohorts, $U$ ($e\notin U$). Researchers should carefully specify the choice set $U$ in practice. If a never exposed cohort $E_{i}=\infty$ exists, we can set $U=\{\infty\}$ and estimate $CLATT_{e,e+l}$ for all cohorts ($e\in\{2,\dots,T\}$) and all relative time periods ($l\geq 0$). If no never exposed cohorts exist, we can set $U=\max\{E_{i}\}$, the last exposed cohort, and can estimate $CLATT_{e,e+l}$ for cohort $e\in\{2,\dots,\max\{E_{i}\}-1\}$ and relative time period $l\in\{0,\dots,\max\{E_{i}\}-1-e\}$ because every unit will be exposed to the instrument after time $t=\max\{E_{i}\}$.
@@
$Y_{i,t}=\beta^{e,l}_{0}+\beta^{e,l}_{i,.}\mathbf{1}\{E_{i}=e\}+\beta^{e,l}_{,.t}\mathbf{1}\{T_{i}=e+l\}+\beta^{e,l}_{IV}D_{i,t}+\epsilon^{e,l}_{i,t},$
@@
Here, $E_{N}[\cdot]$ is the sample analog of the conditional expectation. Note that the IV estimator $\widehat{CLATT}_{e,e+l}$ is the Wald-DID estimator for each $CLATT_{e,e+l}$, where the choice set $U$ is an unexposed group and time $t=e-1$ is a pre-exposed period. From the estimation procedure, we call this a stacked two-stage least squares (STS) estimator. [^18] Our STS estimator is related to the DID estimators in staggered DID designs proposed by Callaway and Sant’Anna (2021) and Sun and Abraham (2021). The DID estimator of the treatment and the outcome in our STS estimator corresponds to that of Sun and Abraham (2021), and coincides with that of Callaway and Sant’Anna (2021) for the case when no covariates exist and never treated units are used as a control group. Their proposed methods avoid the issue of TWFE estimators in staggered DID designs, while our STS estimator avoids the issue of TWFEIV estimators in staggered DID-IV designs.
@@
Suppose Assumptions 8-14 hold. Then, the STS estimator $\widehat{CLATT}_{e,e+l}$ is consistent and asymptotically normal.
@@
Our estimation procedure is the same if we have a non-binary, ordered treatment or use repeated cross section data. Online Appendix Section B presents the influence function of the STS estimator in repeated cross section settings.